OpenStudy (anonymous):

An industrial production process costs C(q) million dollars to produce q million units; these units then sell for R(q) million dollars. If C(2.1)=5.1 , R(2.1)=6.6 , MC(2.1)=0.4 and MR(2.1)=0.5, calculate the following. (a) The profit earned by producing 2.1 million units. (b) The approximate change in revenue if production increases from 2.1 to 2.14 million units. (c) The approximate change in revenue if production decreases from 2.1 to 2.07 million units. (d) The approximate change in profit in parts (b) and (c).

OpenStudy (anonymous):

I am assuming MC and MR means Marginal Cost and Marginal Revenue respectively

OpenStudy (anonymous):

yes it does.

OpenStudy (anonymous):

a)Profit(2.1)= Revenue(2.1)-Cost(2.1) = 6.6-5.1=1.5

OpenStudy (anonymous):

b)Use linearization L(x) =F'(x)dx+F(x) R(2.14)=MR(2.1)(0.04)+R(2.1) 0.4*0.04 + 6.6=6.616

OpenStudy (anonymous):

what is linearization?

OpenStudy (anonymous):

b)Use linearization L(x) =F'(x)dx+F(x) R(2.14)=MR(2.1)(0.04)+R(2.1) 0.4*(-0.03) + 6.6=6.588

OpenStudy (anonymous):

Linearization means you are approximating something when you know its derivative

OpenStudy (anonymous):

I'm not sure I understand the concept, I'm sorry.

OpenStudy (anonymous):

Okay, this is detours, to understand the concept of linearization Say for example, you want to know Square root of 9.2 without calculator,you may use linearization to approximate. f(x)=Sqrt[x] f'(x)= 1/2 x^-1/2 We don't know what square root of 9.2 is but we know square root of 9 is 3 f(9.2) = f(9)+f'(9)(.2) f(9.2) 3 +1/6 (.2)=3.0333

OpenStudy (anonymous):

ohh, okay, I understand. So I will linearize for the whole problem?

OpenStudy (anonymous):

yep, I hope it makes sense

OpenStudy (anonymous):

This is basically we did for b) and c)

OpenStudy (anonymous):

so for c, I subtract instead of add.

OpenStudy (anonymous):

the number we multiply is negative \[0.4*\underline{(-0.03)} + 6.6=6.588\]

OpenStudy (anonymous):

Do you know how I got 0.03 ?

OpenStudy (anonymous):

where did you get that from?

OpenStudy (anonymous):

if production decreases from 2.1 to 2.07 million units. 2.1-2.07=.03 negative because we decrease

OpenStudy (anonymous):

oh oh makes sense, I missed that. Okay so then for the last part, I average b and c together, right?

OpenStudy (anonymous):

but why did we get 6.6 and 6.588?

OpenStudy (anonymous):

so to get jusssssssssssssst the change, i only multiply?so for d, i average them?

OpenStudy (anonymous):

They want the CHANGE in profit. profit=reveune - cost dp= dr-dc You know dr, you don't know dc(change in cost),tying to find that for b)

OpenStudy (anonymous):

I don't understand:x

OpenStudy (anonymous):

Okay, hold on MR dx b) change in revenue= 0.5*0.04= 0.02 c) change in revenue= 0.5*(-0.03)=-0.015 I incorrectly use MC instead of MR above

OpenStudy (anonymous):

Let's do d) We want change in cost for b) MC * dx 0.4*0.04 =0.016 c) 0.4 -.03=-0.012

OpenStudy (anonymous):

Remember Profit= Revenue - Cost To get change in Profit= Change in Rev - Change in Cost for b) 0.02 - 0.016=0.004 for c) -0.015 - -0.012=-0.003

OpenStudy (anonymous):

so 0.02 in b would be what in dollars though?

OpenStudy (anonymous):

change in revenue in b 0.02 million would be $20,000

OpenStudy (anonymous):

I'm trying to understand d.

OpenStudy (anonymous):

What doesn't make sense?

OpenStudy (anonymous):

the change in cost part

OpenStudy (anonymous):

It is same as what we did with revenue but we used MC instead of MR

OpenStudy (anonymous):

oh okay! I think i get it:)

OpenStudy (anonymous):

so .004 would be 4000 dollars right? so hypothetically, .024 would be 24000?

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

You multiply by a million

OpenStudy (anonymous):

I got everything right except d:x

OpenStudy (anonymous):

did you put for 4000, -3000 for d?

OpenStudy (anonymous):

it like reversed values.

OpenStudy (anonymous):

you means it was -3000 and 40000

OpenStudy (anonymous):

and then when I plugged in values for the next problem, I got the same thing, everything right except d.

OpenStudy (anonymous):

Can you check to see if this value for d is right $18400,-13800

OpenStudy (anonymous):

unfortunately if i get the problem wrong, it gives me another:x

OpenStudy (anonymous):

I see, Post you question, I will try to solve it; If I am right, I will explains how I did it

OpenStudy (anonymous):

An industrial production process costs C(q) million dollars to produce q million units; these units then sell for R(q) million dollars. If C(2.1)=5.4 R(2.1)=6.5, MC(2.1)=0.3and MR(2.1)=0.4 calculate the following. (a) The profit earned by producing 2.1 million units. (b) The approximate change in revenue if production increases from 2.1 to 2.12million units. (c) The approximate change in revenue if production decreases from 2.1 to 2.03million units. (d) The approximate change in profit in parts (b) and (c).

OpenStudy (anonymous):

a)1.1 b)0.008 c)-0.028 d) 0.001,-0.008

OpenStudy (anonymous):

i got 2000 and -7000 for d

OpenStudy (anonymous):

Try my answer

OpenStudy (anonymous):

still wrong, but I got the next one right:) you subtract by change in cost then convert to thousand dollars

OpenStudy (anonymous):

thankyou for all your help!:) i appreciate it!